OpenStudy (anonymous):

Who is really smart at math, and that can help me with a few questions? And fast plz.

OpenStudy (anonymous):

what kind?

OpenStudy (anonymous):

like this one: -5/6 - 1/2= -2/3 + -5/6= -1/6 + 2/3=

OpenStudy (anonymous):

@angel12310 do you know?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-5/6 − -1/2=-1/3

OpenStudy (anonymous):

OK THANKS

OpenStudy (anonymous):

-2/3 + -5/6=-1 1/2

OpenStudy (anonymous):

are you working on the other ones?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

-1/6 + 2/3=1/2

OpenStudy (anonymous):

there i did them all

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do you have time to help me with 7 more? sry

OpenStudy (anonymous):

yes i have all the time in the world

OpenStudy (anonymous):

LOL ok thanks do you fo FLVS?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Florida Vitrual School, or are you like in college?

OpenStudy (anonymous):

i'm in middle school in Nashville TN

OpenStudy (anonymous):

i'm homeschooled but give me some more questions

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

here comes the ques

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i'm ready

OpenStudy (anonymous):

7/10n = 1/4 ( what is n) n/2= 4/5 (what is n) n/8= 6/7 (what is n) 1 2/3n = 5/8 (what is n)

OpenStudy (anonymous):

i can't do it with numbers sorry

OpenStudy (anonymous):

these are the choices: 1 3/5, 3/8, 5/14, 6 6/7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so can you do it now?

OpenStudy (anonymous):

none of those are it

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

so do I give you the other questions that I had?

OpenStudy (anonymous):

i can't do it

OpenStudy (anonymous):

only regular fractions ok

OpenStudy (anonymous):

or fractions into decimals

OpenStudy (anonymous):

let me get someone he might help

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

I a confused, what's the q?

OpenStudy (anonymous):

he's here

OpenStudy (anonymous):

ok look

OpenStudy (anonymous):

he needs help with these:

OpenStudy (anonymous):

7/10n = 1/4 ( what is n) n/2= 4/5 (what is n) n/8= 6/7 (what is n) 1 2/3n = 5/8 (what is n)

OpenStudy (anonymous):

@SolomonZelman:7/10n = 1/4 ( what is n) n/2= 4/5 (what is n) n/8= 6/7 (what is n) 1 2/3n = 5/8 (what is n) These are the choices: these are the choices: 1 3/5, 3/8, 5/14, 6 6/7

OpenStudy (anonymous):

i don't get it do you @SolomonZelman

OpenStudy (anonymous):

@angel12310 I'm a girl not a boy..... LOL thats ok

OpenStudy (anonymous):

i know sorry

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

Is @SolomonZelman there?

OpenStudy (solomonzelman):

yes I am,

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

@SolomonZelman do you get the questions?

OpenStudy (solomonzelman):

which one?

OpenStudy (solomonzelman):

there are 5 of them

OpenStudy (anonymous):

7/10n = 1/4 ( what is n) n/2= 4/5 (what is n) n/8= 6/7 (what is n) 1 2/3n = 5/8 (what is n) these are the choices: 1 3/5, 3/8, 5/14, 6 6/7

OpenStudy (anonymous):

Sorry but while @SolomonZelman tries to work then out, @angel12310 can you help me with one that only has fractions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what are the fraction questions

OpenStudy (solomonzelman):

idk what's going on, I can help all of you guys, just ask.

OpenStudy (anonymous):

i don't need help

OpenStudy (anonymous):

maybe tomorrow

OpenStudy (anonymous):

@angel12310 : add 2 1/3 + -4 2/5 = -2 1/15, -2 1/2, -1 14/15, and -6 11/15

OpenStudy (anonymous):

@happyface7 is giving me fraction questions

OpenStudy (anonymous):

what happen?

OpenStudy (anonymous):

do you know how to answer that @angel12310?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you plz answer the question...

OpenStudy (anonymous):

add 2 1/3 + -4 2/5 = -2 1/15, -2 1/2, -1 14/15, and -6 11/15

OpenStudy (anonymous):

7/3 + -22/5=-2 1/15

OpenStudy (anonymous):

so the answer is 21/15?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

improper

OpenStudy (anonymous):

i mean no

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

can u plz help me with the last question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2 1/15

OpenStudy (anonymous):

Subtract 4 1/2 - 3 5/8= Choices: 1 1/8, 7/8, 1/8, and 1 7/8

OpenStudy (anonymous):

do you need help with changing fractions into decimals

OpenStudy (anonymous):

nope just that last question

OpenStudy (anonymous):

= 9/2 − 29/8 = ((9 × 8) − (29 × 2)) / (2 × 8) = (72 - 58) / 16 = 14/16 = 7/8

OpenStudy (anonymous):

so the answer is 7/8

OpenStudy (anonymous):

Subtract 4 1/2 - 3 5/8= Choices: 1 1/8, 7/8, 1/8, and 1 7/8

OpenStudy (anonymous):

oh ok thanks

OpenStudy (anonymous):

anymore of those it's easy

OpenStudy (anonymous):

you really help alot, what grade are you in?

OpenStudy (anonymous):

6th grade i'm 11

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

is that bad

OpenStudy (anonymous):

are you online everyday?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no that is not bad

OpenStudy (anonymous):

except Saturdays and Sundays well sometimes on saturdays

OpenStudy (anonymous):

ok well thank you so much and if i need help with anything else I will try to find you on open study. Thank for being my fan

OpenStudy (anonymous):

welcome just do this to get me:

OpenStudy (anonymous):

@angel12310